# Remote Sensing Image Intelligent Interpretation with the Language-Centered Perspective: Principles, Methods and Challenges

### v1

@@ III-C Language-Centered Knowledge Association
The structure of this section is illustrated in Fig.9, and representative methods for each component are summarized in Table II.
TABLE II: Language-Centered knowledge association structure and representative methods.
@@ III-C1 Task Understanding
The depth of task understanding in LLMs is reflected in their ability to decompose complex tasks. For example, a ship motion monitoring task requires the interpretation system to detect and recognize all ships in a specified maritime region, identify their fine-grained categories, and determine dynamic changes over time through temporal analysis. Due to the highly composite nature of this task, a single model would struggle to handle it effectively. Instead, an LLM is capable of decomposing it into four relatively independent yet sequentially connected subtasks: detection, coarse classification, fine-grained recognition, and temporal analysis each of which can be delegated to a specialized expert model. In this process, the LLM effectively re-structures the complex task of ”ship motion monitoring” into a series of simpler subtasks, each with a clear knowledge objective and a corresponding expert model. This demonstrates the LLM’s comprehension of the task and also reduces the overall task complexity, thereby facilitating more efficient task execution. From this perspective, based on whether the task environment including the knowledge space and the external environment changes during task understanding, we categorize task understanding into two types: Closed-Form task understanding, and Open-Ended task understanding.
@@
In closed-form task understanding scenarios, both the knowledge space and the environment space are static and fixed. Task understanding is primarily achieved through existing knowledge, without involving interaction with the external environment. There are three main approaches to task decomposition in this setting: Rule-based task decomposition, Pretrained language model-based task decomposition, and Chain-of-thought-based task understanding.
@@
  • Rule-Based Task Decomposition: Rule-based task decomposition relies on predefined rules and fixed logical procedures to break down and execute tasks, as illustrated in Fig.10(a). These rules are typically handcrafted for specific task requirements. Once the model receives a task input, it follows the established rules step by step to reason and complete the task. This approach offers strong interpretability, as each reasoning step and decision-making rule is explicit and fixed, allowing the process by which the model derives its output from the input to be clearly traced, understood, and verified. In the field of natural language processing, early rule-based systems such as APLICOT[73] (based on Prolog) used logical rules (e.g., A :- B, C, meaning ”IF B AND C, THEN A”) to connect multiple rule units and perform reasoning in complex tasks. However, this approach suffered from limited context understanding, as it could only build logical relationships based on local information. To address this shortcoming, ACT-R[74] extended APLICOT[73] by introducing a cognitive framework that combines fixed logic rules with memory units, enabling higher-level cognitive processing and improving flexibility and efficiency in task decomposition. Although rule-based methods can effectively support task understanding and decomposition under controllable logical chains, their reliance on explicitly defined rules has become a significant limitation in the era of big data. When dealing with complex tasks, the rigidity of rules often leads to poor performance. As a result, with the rise of language model-based methods, rule-based decomposition has been gradually phased out particularly due to language models’ superior flexibility and adaptability in handling diverse task scenarios.
@@
    In the field of remote sensing, complex tasks often rely on expert prior knowledge, with predefined rules used to decompose tasks into subcomponents, which are then processed step by step in conjunction with interpretation models. For example, in the area of target recognition, RS-DFM[75] addresses 3D object detection under multi-platform observation conditions by decomposing the task into four fixed subtasks: feature extraction, BEV (bird’s-eye view) generation, high- and low-frequency feature decoupling and calibration, and feature fusion, in order to improve detection accuracy and robustness. In multimodal interpretation, Feng et al.[76] (2023) break down cross-modal interpretation into two subtasks: multi-scale spatial feature extraction and multimodal feature fusion. In change detection, Semantic-CD[77] decomposes the task into semantic segmentation and change feature extraction. It uses an Adapted Vision Encoder to extract semantic features from bi-temporal images, followed by a semantic change detection decoder to identify areas of semantic variation. These methods demonstrate that rule-based task decomposition plays an important role in remote sensing interpretation, as it not only simplifies complex tasks but also enhances accuracy and operational stability. However, such approaches are heavily dependent on expert-designed priors, with fixed task decomposition structures, making them poorly suited for open and dynamic environments. Their inability to adapt subtask structures according to changing scenes limits their flexibility and generalizability in complex task settings.
@@
  • Pretrained Language Models Based Task Decomposition: As illustrated in Fig.10(b), task decomposition based on pretrained language models leverages large-scale models such as Llama3[78], GPT-3[17], and DeepSeek-V3[79]. These models are trained on massive text corpora and possess the ability to understand and flexibly handle a wide range of tasks. Task decomposition in this paradigm is typically performed by interpreting task descriptions or instructions provided as input, and generating output based on the model’s pretrained knowledge and contextual understanding. The primary advantage of this approach lies in its flexibility: pretrained language models can handle a wide variety of tasks and adapt to diverse input types without the need for manually designed rules or constraints. Thanks to their strong language understanding capabilities, these models can provide reasonable solutions across different scenarios and show superior adaptability to changing task requirements. However, this flexibility comes at a cost—particularly in terms of interpretability. Since the decision-making process of these models is based on internal parameters and massive training data, the reasoning steps are often opaque and difficult to verify externally. As a result, the outputs may be non-traceable and lack guaranteed correctness.
@@
    In the remote sensing domain, as new tasks continue to emerge—such as Remote Sensing Visual Question Answering (RSVQA), Remote Sensing Visual Captioning (RSVC), and Remote Sensing Visual Grounding (RSVG)—interpretation tasks are increasingly expressed in natural language. Rule-based task decomposition methods struggle to adapt to this flexible task format, prompting researchers to explore how to more efficiently ”understand” remote sensing interpretation tasks. To this end, several studies have introduced LLMs to analyze task requirements. For example, LiT-4-RSVQA[80] uses the text encoder of a VLM to interpret task instructions, while GeoGPT[81] employs an LLM to analyze and decompose interpretation tasks. Moreover, works such as GeoChat[32], SkyEyeGPT[82], and EarthGPT[10] further integrate LLMs into interpretation systems to enhance their understanding and decomposition capabilities for complex tasks. While LLM-based task decomposition offers significantly greater flexibility, allowing the system to automatically adapt to different task demands and intelligently break down complex tasks, the lack of transparency in the reasoning process introduces uncertainty regarding the trustworthiness of the decomposition results. This issue limits the practical applicability of such approaches in remote sensing scenarios that require high reliability.
@@
  • Reasoning-Based Task Decomposition: Reasoning-based task decomposition methods combine the flexibility of language models with the interpretability of rule-based systems by guiding the model to perform step-by-step logical reasoning, as illustrated in Fig.10(c). In this approach, reasoning instructions are provided to the model in the form of prompts, alongside the input query. During the generation of subtasks, the model continuously rethinks and refines its outputs, effectively decomposing the task through a progressive reasoning process. This method encourages the model to self-check at each step, ensuring that task execution is both structured and logically sound. The earliest representative of this approach is the Chain-of-Thought (CoT)[26] prompting method, which uses simple cues like ”Let’s think step by step” to guide the model in generating intermediate reasoning steps and clearly showcasing the decomposition process in the final result. However, the reasoning logic in CoT[26] can be difficult to control, leading to unstable outcomes. To address this, methods such as Self-Reflection[83] and Self-Refine[84] introduce reflection mechanisms, enabling the model to self-evaluate the logic and accuracy of its reasoning, thereby improving reliability. Still, these methods rely on manually crafted prompt templates, and their performance on more complex tasks remains limited. To overcome these challenges, OpenAI-O1[85] and DeepSeek-R1[27] apply reinforcement learning to fine-tune the reasoning capabilities of language models. By integrating self-verification into the training process, these models are able to automatically analyze task steps without relying on external prompts, significantly enhancing their ability to understand and solve complex tasks with greater stability and precision.
@@
    In the remote sensing domain, similar approaches have been explored in Remote Sensing ChatGPT[86], which interprets user instructions in natural language and combines them with domain-specific geoscientific knowledge to decompose complex interpretation tasks into subtasks. These subtasks are then handled sequentially by invoking corresponding expert models. This strategy not only enhances the automation of remote sensing interpretation, but also reduces dependence on domain experts. Furthermore, CEOS[87], developed based on the LangChain[108] framework, serves as an Earth observation analysis platform capable of automatically executing tasks such as vegetation analysis. These works demonstrate the potential of combining logical reasoning and task decomposition to improve the efficiency of remote sensing interpretation. However, these methods often rely on general-purpose LLMs guided by CoT-style prompting, and still suffer from inaccuracies when reasoning over specialized geoscientific knowledge. Due to the complexity of remote sensing imagery and the high level of expertise required in geoscience, existing general LLMs lack sufficient domain-specific knowledge and precise reasoning abilities. Currently, efforts to enhance LLMs’ geoscience reasoning through reinforcement learning-based fine-tuning are still in their early stages.
@@
In summary, while closed-task understanding methods offer certain advantages in task decomposition, they also face inherent limitations. Due to the static nature of the knowledge and environment spaces involved in these methods, models are heavily reliant on predefined rules or pretrained knowledge—which often fall short when confronted with dynamic environments or unseen task requirements. This limitation is particularly evident when dealing with tasks that fall outside the coverage of the model’s internal knowledge base, where LLMs often struggle to deliver satisfactory results. Although rule-based and pretrained language model-based methods can effectively handle well-defined tasks and achieve a high level of automation, their adaptability and flexibility are constrained by the static nature of their knowledge and the lack of interpretability in their reasoning processes. Furthermore, while reasoning-based task decomposition methods attempt to enhance reasoning robustness through chain-of-thought prompting and reflection mechanisms, they still encounter difficulties when facing knowledge gaps or complex task scenarios. As a result, closed-task understanding methods continue to face significant challenges in terms of performance and reliability when applied to dynamic and complex real-world tasks.
@@
In open-ended task understanding scenarios, both the knowledge space and the environment space can be dynamically adjusted according to actual task requirements. To overcome the inherent limitations of closed-task understanding, researchers have begun to explore the integration of external information into the task understanding process. Based on the source of external information, open-task understanding can be further divided into two main pathways: Knowledge base–driven task understanding, and Physical world interaction–driven task understanding.
@@
  • Knowledge Base–Driven Task Decomposition: As illustrated in Fig. 11(a), knowledge base–driven open-task understanding enhances a language model’s comprehension by enabling interaction with external knowledge bases. The model retrieves information from these sources that is similar or relevant to the input query, thereby injecting additional knowledge into the task decomposition process. Representative works include Auto-CoT[88], which retrieves task decomposition examples similar to the input instruction and guides the LLM to imitate these examples for more reasonable task breakdowns. Another example is MemoryBank[89], which leverages Retrieval-Augmented Generation (RAG) to extract task-relevant knowledge from an external database and supports the decomposition process accordingly. Because external knowledge bases can be updated in real time, this method allows LLMs to continuously access the latest information and fill knowledge gaps as needed. However, the relevance and usefulness of the retrieved knowledge still depend heavily on the accuracy of the retrieval algorithm, which determines whether the acquired information truly aligns with the task requirements.
@@
  • Physical World Interaction–Driven Task Decomposition: As illustrated in Fig.11(b), physical world interaction–driven open-task understanding acquires the required information through feedback from interactions with the physical environment, enabling the task decomposition process to unfold progressively. The language model initially generates a set of subtasks based on the input query; as actions associated with these subtasks affect the physical world, feedback from the environment is used to refine and optimize subsequent decomposition steps. A representative work in this direction is ReAct[90], which combines task decomposition with action execution. By directly interacting with the physical world, the model determines the next steps of task decomposition based on both environmental feedback and the results of previous task planning. ReSpAct[91] extends this framework by introducing user interaction feedback, allowing the LLM to seek advice, clarify ambiguities, or understand user preferences through dialogue—thereby updating its task decomposition plan in a more responsive manner. This approach—where task understanding is achieved through direct interaction with the physical environment—enables LLMs to dynamically adjust decomposition strategies according to changes in the environment, thus providing greater adaptability and flexibility. However, a key challenge in this line of research remains: how to design efficient and reliable environment interaction mechanisms.
@@
In the field of remote sensing, open-task understanding is still in its early stages of development. One of the most recent relevant studies is RS-Agent[92]. RS-Agent[92] builds upon a pretrained LLM and implements five specialized agents responsible for reasoning, decision-making, memory, knowledge-enhanced retrieval, and decision-enhanced retrieval, respectively. It also establishes integration channels with conventional remote sensing interpretation tools such as object detection models, superpixel reconstruction models, scene classification models, and change detection models.
@@ III-C2 Knowledge Retrieval
Once the task requirements have been parsed and clarified during the task understanding process, the core challenge shifts to how to extract highly relevant information from internal and external knowledge bases to support subsequent reasoning and decision-making. Two fundamental components of knowledge retrieval are knowledge encoding and knowledge querying. Knowledge encoding focuses on transforming external knowledge into representations that the model can understand and utilize effectively in later stages. Knowledge retrieval, on the other hand, is guided by the specific content of the task requirements and aims to selectively extract information that is highly relevant from both internal and external sources. This ensures the model has access to accurate and timely knowledge, which is critical for effective reasoning and task execution. Through precise knowledge retrieval, the model can compensate for its internal knowledge gaps and leverage existing knowledge bases to further improve its capability and accuracy in handling complex tasks.
@@
External knowledge is typically stored in the form of data, and the process of knowledge encoding essentially involves transforming this multimodal data into vector representations and mapping it into a unified vector space to facilitate subsequent knowledge association and retrieval. Within this space, semantically related knowledge should be located close together, while irrelevant knowledge should be distant. The core objective of this process is to enable the model to efficiently search, access, and integrate relevant knowledge. With the advancement of knowledge retrieval technologies, knowledge encoding has evolved through three key stages: Static encoding, Dynamic encoding, and Multimodal joint encoding.
@@
  • Static Knowledge Encoding: As illustrated in Fig.12(a), static knowledge encoding maps different words into a shared semantic space where semantic relationships between words are preserved—for example, between ”Woman” and ”Queen”, or ”Queen” and ”King”. One of the pioneering methods in this domain is Word2Vec[93], which effectively captures both semantic and syntactic relationships between words. It uses two architectures CBOW (Continuous Bag of Words) and Skip-gram to project high-dimensional word representations into a lower-dimensional vector space, ensuring that semantically similar words are positioned closer together. However, a major limitation of Word2Vec[93] is its inability to capture the dynamic semantic shifts of words across different contexts. To address this, GloVe[94] proposes a matrix factorization approach based on word co-occurrence statistics, allowing it to encode global semantic relationships more effectively. Building on this, FastText[95] refines word representation further by breaking words down into subword units. This enables the model to learn internal word structures, thereby improving its ability to represent low-frequency or out-of-vocabulary words. Nevertheless, static knowledge encoding fails to reflect contextual semantic variation. As shown in Fig.12(a), methods like Word2Vec[93] tend to map the word ”bank”—regardless of whether it refers to a financial institution or a riverbank—to the same point in the semantic space, leading to semantic ambiguity.
@@
  • Dynamic Knowledge Encoding: Unlike static encoding methods, which assign a fixed vector representation to each word, dynamic knowledge encoding generates context-dependent word representations, enabling more accurate semantic interpretation, especially in the presence of contextual variability and polysemy. As illustrated in Fig.12(b), the word ”bank” can be mapped to different semantic regions depending on its surrounding context—whether referring to a financial institution or a riverbank. A representative method in this category is ELMo[96], which utilizes a pretrained bidirectional language model to produce dynamic word embeddings based on surrounding context, effectively capturing the context-specific meanings of words. BERT[71] takes this a step further by employing the bidirectional self-attention mechanism of the Transformer architecture, allowing it to simultaneously consider both the left and right context of a word to generate highly context-aware embeddings. In contrast to BERT[71], GPT[97] is a unidirectional language model trained to predict the next token in a sequence. This training strategy enables GPT[97] to compress semantic information from the input corpus efficiently, despite not incorporating full bidirectional context.
@@
  • Multimodal Joint Encoding: While dynamic knowledge encoding has made significant progress in handling language-based knowledge representation, it struggles to effectively process multimodal data. In contrast, multimodal knowledge encoding offers a more comprehensive and unified solution for integrating information across diverse modalities. As illustrated in Fig.12(c), this approach enables the joint representation of vision, language, and other modalities to describe a shared semantic concept, laying a solid foundation for downstream knowledge retrieval and reasoning. In this field, MKBE[98] constructs a multimodal knowledge base by unifying the representations of visual and language modalities. During the construction process, MKBE[98] focuses on preserving temporal characteristics of each modality and the relationships between entities expressed across modalities. For example: For the visual modality, MKBE[98] uses Convolutional Neural Networks (CNNs) to extract spatial and semantic features from images. For the language modality, LSTM[109] networks are employed to capture sequential dependencies between words. At the entity-relation level, MKBE[98] explicitly expresses cross-modal relationships by designing subject-relation pairs as training ground truth, and uses a cross-entropy loss function to train the model to recognize entity relationships. Building on MKBE[98], MULE[99] further enriches the language modality by supporting multilingual representations, thereby improving the model’s compatibility with multiple languages. Although these methods have made notable advances in multimodal joint encoding, they still require dedicated embedding modules for each individual modality. With the rapid development of LLMs, newer approaches like E5-V[100] leverage MLLMs as unified encoders, allowing different modalities to be directly mapped into a shared embedding space without the need for modality-specific components. This approach significantly simplifies the implementation of multimodal joint encoding and improves the efficiency of cross-modal data integration.
@@
    In the field of remote sensing, the unified representation and understanding of multimodal data remains a central challenge for intelligent interpretation systems. The main difficulty lies in effectively fusing information from different modalities. With the rapid advancement of LLMs, a language-centered paradigm for multimodal joint embedding has gained increasing attention. EarthMarker[11] applies this paradigm to remote sensing by embedding visual modality data into the representational space of language models, thereby constructing a modality-unified semantic space. Within this framework, EarthMarker[11] introduces the MoV (Mixture of Visual Experts) and a modality alignment module to integrate remote sensing imagery, visual object prompts, and textual instructions. This enables unified understanding and representation of heterogeneous multimodal information. This approach not only enhances the interpretability of remote sensing data, but also offers a novel pathway for intelligent processing in multimodal tasks.

### v2

@@ III-C Language-Centered Knowledge Association and Reasoning
The structure of this section is illustrated in Fig.9, and representative methods for each component are summarized in Table III.
TABLE III: Language-centered knowledge association structure and representative methods.
@@ III-C1 Task Understanding
The depth of task understanding in LLMs is reflected in their ability to decompose complex tasks. For example, a ship motion monitoring task requires an interpretation system to detect and recognize all ships in a specified maritime region, identify their fine-grained categories, and determine dynamic changes over time through temporal analysis. Owing to the highly composite nature of this task, a single model would struggle to handle it effectively. Instead, an LLM is capable of decomposing it into four relatively independent yet sequentially connected subtasks, namely, detection, coarse classification, fine-grained recognition, and temporal analysis, each of which can be delegated to a specialized expert model. In this process, the LLM effectively restructures the complex task of “ship motion monitoring” into a series of simpler subtasks, each with a clear knowledge objective and a corresponding expert model. This operation demonstrates the LLM’s comprehension of the task and reduces the overall task complexity, thereby facilitating more efficient task execution. From this perspective, based on whether the task environment, including the knowledge space and the external environment, changes during task understanding, we categorize task understanding into two types: closed-form task understanding and open-ended task understanding.
@@
In closed-form task understanding scenarios, both the knowledge space and the environment space are static and fixed. Task understanding is achieved primarily through existing knowledge, without involving interaction with the external environment. There are three main approaches to task decomposition in this setting: rule-based task decomposition, pretrained language model-based task decomposition, and chain-of-thought-based task understanding.
@@
  • Rule-Based Task Decomposition: Rule-based task decomposition relies on predefined rules and fixed logical procedures to break down and execute tasks, as illustrated in Fig.10(a). These rules are typically handcrafted for specific task requirements. Once the model receives a task input, it follows the established rules step by step to reason and complete the task. This approach offers strong interpretability, as each reasoning step and decision-making rule is explicit and fixed, allowing the process by which the model derives its output from the input to be clearly traced, understood, and verified. In the field of natural language processing, early rule-based systems such as APLICOT[117] (based on Prolog) used logical rules (e.g., A :- B, C, meaning “IF B AND C, THEN A”) to connect multiple rule units and perform reasoning in complex tasks. However, this approach suffered from limited context understanding, as it could only build logical relationships based on local information. To address this shortcoming, ACT-R[8] extended APLICOT[117] by introducing a cognitive framework that combines fixed logic rules with memory units, enabling higher-level cognitive processing and improving flexibility and efficiency in task decomposition. Although rule-based methods can effectively support task understanding and decomposition under controllable logical chains, their reliance on explicitly defined rules has become a significant limitation in the era of big data. When dealing with complex tasks, the rigidity of rules often leads to poor performance. As a result, with the rise of language model-based methods, rule-based decomposition has been gradually phased out, particularly because of language models’ superior flexibility and adaptability in handling diverse task scenarios.
@@
    In the field of remote sensing, complex tasks often rely on expert prior knowledge, with predefined rules used to decompose tasks into subcomponents, which are then processed step by step in conjunction with interpretation models. For example, in the area of target recognition, RS-DFM[171] addresses 3D object detection under multiplatform observation conditions by decomposing the task into four fixed subtasks to improve detection accuracy and robustness: feature extraction, BEV (bird’s-eye view) generation, high- and low-frequency feature decoupling and calibration, and feature fusion. In multimodal interpretation, Feng et al.[38] (2023) break down cross-modal interpretation into two subtasks: multiscale spatial feature extraction and multimodal feature fusion. In change detection, semantic-CD[209] decomposes the task into semantic segmentation and change feature extraction. It uses an adapted vision encoder to extract semantic features from bitemporal images, followed by a semantic change detection decoder to identify areas of semantic variation. These methods demonstrate that rule-based task decomposition plays an important role in remote sensing interpretation, as it not only simplifies complex tasks but also enhances accuracy and operational stability. However, such approaches are heavily dependent on expert-designed priors, with fixed task decomposition structures, making them poorly suited for open and dynamic environments.
@@
  • Pretrained Language Model-based Task Decomposition: As illustrated in Fig.10(b), task decomposition based on pretrained language models leverages large-scale models such as Llama3[40], GPT-3[17], and DeepSeek-V3[91]. These models are trained on massive text corpora and can understand and flexibly handle a wide range of tasks. Task decomposition in this paradigm is typically performed by interpreting task descriptions or instructions provided as input and generating output based on the model’s pretrained knowledge and contextual understanding. The primary advantage of this approach lies in its flexibility: pretrained language models can handle a wide variety of tasks and adapt to diverse input types without the need for manually designed rules or constraints. Owing to their strong language understanding capabilities, these models can provide reasonable solutions across different scenarios and show superior adaptability to changing task requirements. However, this flexibility comes at a cost–particularly in terms of interpretability. Since the decision-making process of these models is based on internal parameters and massive training data, the reasoning steps are often opaque and difficult to verify externally. As a result, the outputs may be nontraceable and lack guaranteed correctness.
@@
    In the remote sensing domain, as new tasks continue to emerge—such as remote sensing visual question answering (RSVQA), remote sensing visual captioning (RSVC), and remote sensing visual grounding (RSVG)—interpretation tasks are being increasingly expressed in natural language. Rule-based task decomposition methods struggle to adapt to this flexible task format, prompting researchers to explore how to more efficiently “understand” remote sensing interpretation tasks. To this end, several studies have introduced LLMs to analyze task requirements. For example, LiT-4-RSVQA[44] uses the text encoder of a VLM to interpret task instructions, while GeoGPT[201] employs an LLM to analyze and decompose interpretation tasks. Moreover, works such as GeoChat[74], SkyEyeGPT[196], and EarthGPT[199] further integrate LLMs into interpretation systems to enhance their understanding and decomposition capabilities for complex tasks. While LLM-based task decomposition offers significantly greater flexibility, allowing the system to automatically adapt to different task demands and intelligently break down complex tasks, the lack of transparency in the reasoning process introduces uncertainty regarding the trustworthiness of the decomposition results. This issue limits the practical applicability of such approaches in remote sensing scenarios that require high reliability.
@@
  • Reasoning-Based Task Decomposition: Reasoning-based task decomposition methods combine the flexibility of language models with the interpretability of rule-based systems by guiding the model to perform step-by-step logical reasoning, as illustrated in Fig.10(c). In this approach, reasoning instructions are provided to the model in the form of prompts, alongside the input query. During the generation of subtasks, the model continuously rethinks and refines its outputs, effectively decomposing the task through a progressive reasoning process. This method encourages the model to self-check at each step, ensuring that task execution is both structured and logically sound. The earliest representative of this approach is the chain-of-thought (CoT)[176] prompting method, which uses simple cues like “Let’s think step by step” to guide the model in generating intermediate reasoning steps and clearly showcasing the decomposition process in the final result. However, the reasoning logic in CoT[176] can be difficult to control, leading to unstable outcomes. To address this issue, methods such as self-reflection[136] and self-refinement[110] introduce reflection mechanisms, enabling the model to self-evaluate the logic and accuracy of its reasoning, thereby improving reliability. Nevertheless, these methods rely on manually crafted prompt templates, and their performance on more complex tasks remains limited. To overcome these challenges, OpenAI-O1[60] and DeepSeek-R1[42] apply reinforcement learning to fine-tune the reasoning capabilities of language models.
@@
    In the remote sensing domain, similar approaches have been explored in remote sensing ChatGPT[43], which interprets user instructions in natural language and combines them with domain-specific geoscientific knowledge to decompose complex interpretation tasks into subtasks. These subtasks are then handled sequentially by invoking corresponding expert models. This strategy not only enhances the automation of remote sensing interpretation but also reduces the dependence on domain experts. Furthermore, CEOS[165], which was developed based on the LangChain[124] framework, serves as an Earth observation analysis platform capable of automatically executing tasks such as vegetation analysis.
@@
In summary, while closed-task understanding methods offer certain advantages in terms of task decomposition, they also face inherent limitations. Owing to the static nature of the knowledge and environment spaces involved in these methods, models are heavily reliant on predefined rules or pretrained knowledge, which often fall short when confronted with dynamic environments or unseen task requirements. This limitation is particularly evident when dealing with tasks that fall outside the coverage of the model’s internal knowledge base, where LLMs often struggle to deliver satisfactory results. Although rule-based and pretrained language model-based methods can effectively handle well-defined tasks and achieve a high level of automation, their adaptability and flexibility are constrained by the static nature of their knowledge and the lack of interpretability in their reasoning processes. Furthermore, while reasoning-based task decomposition methods attempt to improve reasoning robustness through chain-of-thought prompting and reflection mechanisms, they still encounter difficulties when facing knowledge gaps or complex task scenarios. As a result, closed-task understanding methods continue to face significant challenges in terms of performance and reliability when applied to dynamic and complex real-world tasks.
@@
In open-ended task understanding scenarios, both the knowledge space and the environment space can be dynamically adjusted according to actual task requirements. To overcome the inherent limitations of closed-task understanding, researchers have begun to explore the integration of external information into the task understanding process. Based on the source of external information, open-task understanding can be further divided into two main pathways: knowledge base-driven task understanding and physical world interaction-driven task understanding.
@@
  • Knowledge Base-Driven Task Decomposition: As illustrated in Fig. 11(a), knowledge base-driven open-task understanding enhances a language model’s comprehension by enabling interaction with external knowledge bases. The model retrieves information from these sources that is similar or relevant to the input query, thereby injecting additional knowledge into the task decomposition process. Representative works include Auto-CoT[203], which retrieves task decomposition examples similar to the input instruction and guides the LLM to imitate these examples for more reasonable task breakdowns. Another example is MemoryBank[205], which leverages retrieval-augmented generation (RAG) to extract task-relevant knowledge from an external database and supports the decomposition process accordingly. Because external knowledge bases can be updated in real time, this method allows LLMs to continuously access the latest information and fill knowledge gaps as needed. However, the relevance and usefulness of the retrieved knowledge still depend heavily on the accuracy of the retrieval algorithm, which determines whether the acquired information truly aligns with the task requirements.
@@
  • Physical World Interaction-Driven Task Decomposition: As illustrated in Fig.11(b), physical world interaction-driven open-task understanding acquires the required information through feedback from interactions with the physical environment, enabling the task decomposition process to unfold progressively. The language model initially generates a set of subtasks based on the input query; as actions associated with these subtasks affect the physical world, feedback from the environment is used to refine and optimize the subsequent decomposition steps. A representative work in this direction is ReAct[190], which combines task decomposition with action execution. By directly interacting with the physical world, the model determines the next steps of task decomposition based on both environmental feedback and the results of previous task planning. ReSpAct[34] extends this framework by introducing user interaction feedback, allowing the LLM to seek advice, clarify ambiguities, or understand user preferences through dialog, thereby updating its task decomposition plan in a more responsive manner. However, a key challenge in this line of research remains how to design efficient and reliable environmental interaction mechanisms.
@@
In the field of remote sensing, open-task understanding is still in its early stages of development. One of the most recent relevant studies is RS-Agent[184]. RS-Agent[184] builds upon a pretrained LLM and implements five specialized agents that are individually responsible for reasoning, decision-making, memory, knowledge-enhanced retrieval, and decision-enhanced retrieval. It also establishes integration channels with conventional remote sensing interpretation tools such as object detection models, superpixel reconstruction models, scene classification models, and change detection models.
@@ III-C2 Knowledge Retrieval
Once the task requirements have been parsed and clarified during the task understanding process, the core challenge shifts to how to extract highly relevant information from internal and external knowledge bases to support subsequent reasoning and decision-making. Two fundamental components of knowledge retrieval are knowledge encoding and knowledge querying. Knowledge encoding focuses on transforming external knowledge into representations that the model can understand and utilize effectively in later stages. Knowledge retrieval, on the other hand, is guided by the specific content of the task requirements and aims to selectively extract information that is highly relevant from both internal and external sources. This component ensures that the model has access to accurate and timely knowledge, which is critical for effective reasoning and task execution. Through precise knowledge retrieval, the model can compensate for its internal knowledge gaps and leverage existing knowledge bases to further improve its ability and accuracy in handling complex tasks.
@@
External knowledge is typically stored in the form of data. The process of knowledge encoding essentially involves transforming this multimodal data into vector representations and mapping it into a unified vector space to facilitate subsequent knowledge association and retrieval. Within this space, semantically related knowledge should be located close together, while irrelevant knowledge should be distant. The core objective of this process is to enable the model to efficiently search, access, and integrate relevant knowledge. With the advancement of knowledge retrieval technologies, knowledge encoding has evolved through three key stages: static encoding, dynamic encoding, and multimodal joint encoding.
@@
  • Static Knowledge Encoding: As illustrated in Fig.12(a), static knowledge encoding maps different words into a shared semantic space where semantic relationships between words are preserved—for example, between “Woman” and “Queen”, or “Queen” and “King”. One of the pioneering methods in this domain is word2vec[29], which effectively captures both semantic and syntactic relationships between words. It uses two architectures, CBOW (continuous bag of words) and skip-gram, to project high-dimensional word representations into a lower-dimensional vector space, ensuring that semantically similar words are positioned closer together. However, a major limitation of word2vec[29] is its inability to capture the dynamic semantic shifts of words across different contexts. To address this shortcoming, GloVe[128] proposed a matrix factorization approach based on word co-occurrence statistics, allowing it to encode global semantic relationships more effectively. Building on this work, FastText[67] refined word representation further by breaking words down into subword units. This capability enables the model to learn internal word structures, thereby improving its ability to represent low-frequency or out-of-vocabulary words. Nevertheless, static knowledge encoding fails to reflect contextual semantic variation. As shown in Fig.12(a), methods such as word2vec[29] tend to map the word “bank”—regardless of whether it refers to a financial institution or a riverbank—to the same point in the semantic space, leading to semantic ambiguity.
@@
  • Dynamic Knowledge Encoding: Unlike static encoding methods, which assign a fixed vector representation to each word, dynamic knowledge encoding generates context-dependent word representations, enabling more accurate semantic interpretation, especially in the presence of contextual variability and polysemy. As illustrated in Fig.12(b), the word “bank” can be mapped to different semantic regions depending on its surrounding context-’whether it refers to a financial institution or a riverbank. A representative method in this category is ELMo[138], which uses a pretrained bidirectional language model to produce dynamic word embeddings based on the surrounding context, effectively capturing the context-specific meanings of words. BERT[72] takes this approach a step further by employing the bidirectional self-attention mechanism of the transformer architecture, thus allowing it to simultaneously consider both the left and right context of a word to generate highly context-aware embeddings. In contrast to BERT[72], GPT[135] is a unidirectional language model trained to predict the next token in a sequence. This training strategy enables GPT[135] to compress semantic information from the input corpus efficiently, despite not incorporating the full bidirectional context.
@@
  • Multimodal Joint Encoding: While dynamic knowledge encoding has made significant progress in handling language-based knowledge representation, it struggles to effectively process multimodal data. In contrast, multimodal knowledge encoding offers a more comprehensive and unified solution for integrating information across diverse modalities. As illustrated in Fig.12(c), this approach enables the joint representation of vision, language, and other modalities to describe a shared semantic concept, laying a solid foundation for downstream knowledge retrieval and reasoning. In this field, MKBE[129] constructs a multimodal knowledge base by unifying the representations of visual and language modalities. During the construction process, MKBE[129] focuses on preserving the temporal characteristics of each modality and the relationships between entities expressed across modalities. For example, for the visual modality, MKBE[129] uses convolutional neural networks (CNNs) to extract spatial and semantic features from images. For the language modality, LSTM[47] networks are employed to capture sequential dependencies between words. At the entity-relation level, MKBE[129] explicitly expresses cross-modal relationships by designing subject-relation pairs as training ground truth and uses a cross-entropy loss function to train the model to recognize entity relationships. Building on MKBE[129], MULE[69] further enriches the language modality by supporting multilingual representations, thereby improving the model’s compatibility with multiple languages. Although these methods have made notable advances in multimodal joint encoding, they still require dedicated embedding modules for each individual modality. With the rapid development of LLMs, newer approaches such as E5-V[63] leverage MLLMs as unified encoders, allowing different modalities to be directly mapped into a shared embedding space without the need for modality-specific components. This approach significantly simplifies the implementation of multimodal joint encoding and improves the efficiency of cross-modal data integration.
@@
    In the field of remote sensing, the unified representation and understanding of multimodal data remains a central challenge for intelligent interpretation systems. The main difficulty lies in effectively fusing information from different modalities. With the rapid advancement of LLMs, a language-centered paradigm for multimodal joint embedding has gained increasing attention. EarthMarker[198] applies this paradigm to remote sensing by embedding visual modality data into the representational space of language models, thereby constructing a modality-unified semantic space. Within this framework, EarthMarker[198] introduces MoV (Mixture of Visual Experts) and a modality alignment module to integrate remote sensing imagery, visual object prompts, and textual instructions. This combination enables a unified understanding and representation of heterogeneous multimodal information. This approach not only enhances the interpretability of remote sensing data but also offers a novel pathway for intelligent processing in multimodal tasks.